# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v2

@@ 2.2.2 Hierarchical Normalization
Hydrological magnitudes vary by six orders across sites (0.1 to 100,000 ft3/s), creating severe optimization challenges. Standard approaches fail: global standardization yields large-river dominance; site-specific alone loses cross-site comparability; min-max scaling suffers outlier sensitivity; raw units prevent convergence.
We introduce a three-tier hierarchical normalization ( full mathematics in Appendix A) simultaneously achieving: (1) stable gradients across extreme heterogeneity, (2) zero train-test leakage, (3) exact physical unit recovery, and (4) scale-dependent information preservation.
Tier 1 - Logarithmic stabilization: Apply $Q^{(1)}=\ln(Q+\epsilon)$, $H^{(1)}=\ln(H+\epsilon)$ for approximately log-normal variables [45], linearizing rating curves and stabilizing variance.
Tier 2 - Site-specific standardization: Compute $\mu_{s}$, $\sigma_{s}$ exclusively from each site’s training data, then transform $\mathbf{x}^{(2)}=(\mathbf{x}^{(1)}-\mu_{s})/(\sigma_{s}+\epsilon)$. For validation/test sites, substitute global training statistics, maintaining strict partition separation.
Tier 3 - Global clipping: Apply $\mathbf{x}^{\text{norm}}=\text{clip}(\mathbf{x}^{(2)},-3,+3)$ to inputs only, retaining 99.7% of normal variation while preventing gradient explosion. Outputs denormalize without clipping.
Exact inverse: $\hat{y}=\exp[\hat{y}^{\text{norm}}\cdot(\sigma_{s}+\epsilon)+\mu_{s}]-\epsilon$ recovers physical units. Scale embeddings ($\sigma_{\ln Q}$, $\sigma_{\ln H}$) return absolute variability information enabling scale-dependent behavior (e.g., flashiness in small basins).
Rationale: The logarithm converts multiplicative noise to additive; site-specific standardization enables weight sharing without large-river dominance; clipping provides numerical stability; embeddings distinguish 10 ft3/s (small creek) from 10,000 ft3/s (large river) despite identical post-normalization values. No single-tier normalization achieves all requirements for continental-scale learning.
@@ 2.3 Training-Time Anomaly Injection
Anomaly detection requires labeled corruption patterns, but manual annotation of 6M sequences is prohibitively expensive. We implement on-the-fly synthetic injection generating diverse patterns during stage 2 fine tuning for anomaly detection which is also a part of our training process.
Deliberate simplification philosophy: The training injector implements $\sim$11 simplified patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing) applied in normalized log-space, deliberately less sophisticated than test anomalies. This design forces learning of fundamental hydrometric consistency principles (discharge-stage coupling, temporal smoothness, physical plausibility) rather than memorizing specific signatures. If training and test used identical patterns, strong performance could arise from pattern matching rather than genuine understanding. The training-test complexity gap creates a defendable generalization test.
Controlled coverage: Two-tier system assigns light corruption (60% probability, 5-15% coverage) or moderate (40% probability, 15-30% coverage) with iterative refinement ($\pm 3\%$ tolerance for light, $\pm 5\%$ for moderate), maintaining mean 15.2% $\pm$ 3.1% across batches. Mixing: single-type (60%) teaches type-specific patterns; double-type (40%) teaches discrimination. Segments: $n\in[2,4]$ with lengths $L\in[T/100,T/4]$ randomly positioned. Curriculum scheduling ramps injection probability from 0.2 (epochs 1-2) to 0.4 (thereafter). Complete implementation details in Appendix E.
2.4 Evaluation Datasets
@@ 2.4.1 Synthetic Test Set Design
Rigorous generalization assessment requires test data preventing memorization while probing genuine abstraction. We construct a synthetic test set from 799 held-out USGS sites with expert-validated anomalies spanning 18 consolidated types (Table 5). Four orthogonal separation axes between training and test distributions ensure strong performance requires learning fundamental hydrometric principles rather than pattern matching.
@@
Mathematical Linear drift, basic offsets Exponential/sigmoid/polynomial; 3-4 variants per type Require functional abstraction
Temporal 8-96 hour segments Micro (3-58h), Meso (7-192h), Macro (72-520h) Probe scale-invariant detection
Parameter Coverage 10-32%, nominal strength Coverage trimodal: 3-9% / 32-44% / 44-60%; Severity bimodal Test distribution boundaries
Geographic separation: Zero site overlap eliminates memorization of station-specific patterns. Mathematical separation: Each test anomaly type implements 3-4 equation variants (e.g., drift: linear, exponential, sigmoid, polynomial) producing similar visual patterns through different mechanisms. Temporal separation: Structured duration regimes (micro/meso/macro) spanning 3-520 hours, with macro extended to 520h based on expert feedback regarding ice persistence [46]. Parameter separation: Coverage and severity anti-overlap the training sweet spot, testing distribution boundaries.
Expert validation and taxonomy consolidation.
We assembled an HTML validation gallery with 8+ representative sequences per anomaly type, reviewed by domain experts using structured protocols (realism scoring, temporal validation, confound detection). Reviewers included a geoscience faculty member with extensive experience in water resources research and a research assistant from the Water Resources Institute at the University of Vermont with operational USGS internship experience examining real-world data quality issues. Based on systematic feedback, we consolidated to 18 types with field-relevance agreement (Table 5), categorized by confidence: high (9 types: dropout, flatline, spike, backwater, ice_backwater, drift, rating_drift, debris_effect, sedimentation), medium (4 types), low (4 types), and debated (retained with caveat). Complete consolidation mapping and expert feedback integration details appear in Appendix G.
@@
Category Types Confidence Variants
Sensor failures Dropout, flatline, spike High 3-4 each
Hydraulic phenomena Backwater, ice_backwater, debris_effect, sedimentation High 3-4 each
Gradual degradation Drift, rating_drift, sensor_fouling High/Medium 3-4 each
Processing errors Bias_step, time_misalignment, quantization Medium/Low 3-4 each
Complex artifacts Splice, warp, noise_burst, gate_operation Low/Debated 3-4 each
@@
All test anomalies inject in denormalized physical space to respect actual discharge-stage coupling. The three-step process: (1) denormalize to physical units, (2) apply parametric transformations with hydraulic constraints (e.g., ice: $H^{\prime}=H(1+\alpha_{\text{ice}})$, $Q^{\prime}=Q(1-\beta_{\text{ice}})$ with $\alpha\sim U(0.15,0.55)$, $\beta\sim U(0,0.10)$), (3) renormalize with clipping. Single-type isolation (30% of sequences) enables per-type diagnostics; compound anomalies (70%, with 40% overlap probability) test discrimination. Complete injection implementation details, equation variants, and reproducibility protocols appear in Appendix F.
Rationale for synthetic evaluation.
Synthetic data provides defendable evaluation: deliberate separation prevents memorization; physical grounding ensures realism; multi-axis design probes distinct generalization modes; expert validation provides external validity; controlled conditions enable fine-grained diagnostics. Complementarity with operational validation (Section 2.4.2) provides stronger evidence than either alone—synthetic proves learned concepts, operational proves practical transfer.
@@ 2.4.2 Canadian Zero-Shot Transfer Evaluation
The USGS archive provides quality-controlled products but lacks systematic raw-vs-corrected record pairs exposing operational correction workflows. To test real-world transfer, we evaluate HydroGEM on Environment and Climate Change Canada (ECCC) stations distributing hourly raw and corrected unit values for both discharge and stage.
@@
Zero-shot protocol: HydroGEM was trained exclusively on USGS sites and never fine-tuned on Canadian data. All experiments are strictly zero-shot, testing whether learned representations generalize across political boundaries, agencies, instrumentation protocols, and rating curve derivation methods.
[…]
@@
Quality filtering: To ensure meaningful evaluation, we applied hydrologically-motivated checks to corrected records: sufficient variability (CV >0.10>0.10), monotonic rating ($\rho>0.5$), reasonable exponent ($b\in[0.5,10]$), moderate fit ($R^{2}\geq 0.3$), valid data fraction ($\geq 70\%$), limited flatlines ($30\%$ differences <0.001<0.001 ft). Station-level filtering excluded sites losing $5\%$ timesteps. Detailed filtering protocols appear in Appendix H.
@@
Weak label construction: We compute relative changes induced by operational correction: $\Delta Q_{\text{rel}}=|Q_{\text{corr}}-Q_{\text{raw}}|/(|Q_{\text{raw}}|+\epsilon)$, similarly for $H$. Timesteps with $\Delta Q_{\text{rel}}>0.01$ or $\Delta H_{\text{rel}}>0.01$ are marked anomalous ($y_{\text{corr}}=1$), providing weak labels encoding where human expertise deemed raw data untrustworthy. The 1% threshold is intentionally conservative, capturing meaningful hydrograph adjustments.
@@
These labels are “weak” because human editors may miss subtle anomalies or correct for operational reasons orthogonal to data quality (e.g., re-rating after channel changes). Nevertheless, they provide externally-validated ground truth unavailable in synthetic datasets. Pattern-based labels (local correlation deviations) complement correction-based labels, emphasizing multi-point segments where corrected hydrograph shape deviates from raw.
@@
Site sampling: From all stations satisfying quality criteria, we randomly sampled 100 distinct stations for evaluation, ensuring representative out-of-sample behavior without cherry-picking. Window filtering required 10-40% correction fraction, identifying substantial but not overwhelming editing effort. For each site, we evaluate HydroGEM in pure zero-shot configuration, comparing binary anomaly masks against correction-based and pattern-based labels.
2.5 Summary
Our data preparation pipeline yields:
  • Foundation corpus: 6.67M sequences (3.84B hourly timesteps) from 5,008 USGS stations (70/15/15 train/val/test split)
  • Training anomalies: On-the-fly injection of $\sim$11 simplified types, 15.2% mean coverage, normalized-space corruption
  • Test anomalies: 18 expert-validated types, 4-axis separation, physical-space injection, 3-4 equation variants per type
  • Real-world transfer: 100 randomly sampled Canadian stations, weak labels from operational corrections, strict zero-shot evaluation
This multi-tiered strategy allows rigorous capability assessment: controlled synthetic tests probe specific generalization modes under known ground truth; real-world Canadian data validates operational transfer. Together, these provide evidence of learned hydrometric consistency principles rather than dataset artifacts.
@@ 3 HydroGEM Model Architecture
Having established the data preparation pipeline in Section 3, we now describe the HydroGEM model that processes the hierarchically normalized 12-dimensional feature vectors to detect and correct anomalies in streamflow time series (Figure 2). The model input $\mathbf{X}\in\mathbb{R}^{T\times 12}$ consists of the features defined in Section 3.2.1, transformed through the three-tier normalization scheme (Section 3.2.2), and segmented into 576-hour windows (Section 3.2.3). Each sequence represents a continental-scale sample spanning diverse hydrologic regimes, requiring the architecture to learn scale-invariant patterns while respecting physical constraints.
Figure 2: HydroGEM training and inference pipeline. Stage 1 pretrains a hybrid TCN-Transformer backbone (14.2M parameters) using masked reconstruction on clean USGS data. Stage 2 freezes the backbone and trains a lightweight detection head ( 11K parameters) on synthetically corrupted sequences. At inference, the model provides anomaly probabilities with uncertainty estimates via MC Dropout, and suggested corrections for human review.

### v3

@@ 2.2.2 Hierarchical Normalization
Hydrological magnitudes vary by 6 orders across sites (0.1 to 100,000 ft3/s), creating severe optimization challenges. Standard approaches are insufficient: global standardization yields large-river dominance; site-specific standardization alone loses cross-site comparability; min-max scaling is sensitive to outliers; and raw units often prevent stable training.
We introduce a 3-tier hierarchical normalization (full mathematics in Appendix A) that achieves (1) stable gradients across extreme heterogeneity, (2) no train-test leakage, (3) exact physical unit recovery, and (4) scale-dependent information preservation.
Tier 1 Logarithmic stabilization: Apply $Q^{(1)}=\ln(Q+\epsilon)$ and $H^{(1)}=\ln(H+\epsilon)$ for approximately log-normal variables [60], which linearizes rating-curve-like relationships and stabilizes variance.
Tier 2 Site-specific standardization: Compute $\mu_{s}$ and $\sigma_{s}$ exclusively from each training site series, then transform $\mathbf{x}^{(2)}=(\mathbf{x}^{(1)}-\mu_{s})/(\sigma_{s}+\epsilon)$. For validation and test sites, we apply global training statistics to avoid leakage and to reflect a deployable setting without site-specific calibration.
Tier 3 Global clipping: Apply $\mathbf{x}^{\text{norm}}=\text{clip}(\mathbf{x}^{(2)},-3,+3)$ to inputs only, retaining 99.7% of typical variation while preventing gradient explosion. Outputs are denormalized without clipping.
Exact inverse: $\hat{y}=\exp[\hat{y}^{\text{norm}}\cdot(\sigma_{s}+\epsilon)+\mu_{s}]-\epsilon$ recovers physical units. Scale embeddings ($\sigma_{\ln Q}$ and $\sigma_{\ln H}$) return absolute variability information, enabling scale-dependent behavior (eg flashiness in small basins).
Rationale: The logarithm converts multiplicative noise to additive; site-specific standardization enables weight sharing without large-river dominance; clipping provides numerical stability; and embeddings distinguish 10 ft3/s from 10,000 ft3/s despite identical normalized values. No single-tier normalization achieves all requirements for continental-scale learning.
@@ 2.3 Training-Time Anomaly Injection
Anomaly detection requires labeled corruption patterns, but manual annotation of millions of sequences is prohibitively expensive. We implement on-the-fly synthetic injection that generates diverse patterns during Stage 2 fine-tuning for anomaly detection.
Deliberate simplification philosophy: The training injector implements approximately 11 simplified patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing) applied in normalized log-space, deliberately less sophisticated than the test anomalies. This design forces learning of fundamental hydrometric consistency principles (discharge-stage coupling, temporal smoothness, physical plausibility) rather than memorizing specific signatures. If training and test used identical patterns, strong performance could arise from pattern matching rather than genuine understanding. The training-test complexity gap creates a defensible generalization test.
Controlled coverage: A 2-tier system assigns light corruption (60% probability, 5-to-15% coverage) or moderate corruption (40% probability, 15-to-30% coverage) with iterative refinement ($\pm 3\%$ tolerance for light, $\pm 5\%$ for moderate), maintaining mean 15.2% $\pm$ 3.1% across batches. Mixing uses single-type corruption (60%) to teach type-specific signatures and double-type corruption (40%) to teach discrimination. Each window uses $n\in[2,4]$ segments with lengths $L\in[T/100,T/4]$ at random positions. Curriculum scheduling ramps injection probability from 0.2 (epochs 1 to 2) to 0.4 thereafter. Complete implementation details appear in Appendix E.
@@ 2.3.1 Synthetic Test Set Design
Rigorous generalization assessment requires test data that prevents memorization and emphasizes functional abstraction. We construct a synthetic test set from 799 held-out USGS sites with anomalies spanning 18 types grounded in USGS operational documentation [59, 31, 30, 62] and peer-reviewed sensor anomaly studies [61, 44, 45]. Four orthogonal separation axes between training and test (Table 5) ensure strong performance requires learning fundamental hydrometric principles rather than pattern matching.
Rationale for synthetic evaluation.
The use of synthetic anomaly injection for benchmarking is established practice in time series anomaly detection research (Table 4). Synthetic injection addresses the scarcity of labeled anomalies in operational archives, enables controlled variation of anomaly characteristics for systematic evaluation, and supports reproducible benchmarking [56, 64]. Our approach differs from prior benchmarks in 2 key ways: (1) anomalies are injected into real USGS hydrological time series rather than fully synthetic data, preserving authentic flow dynamics, and (2) injection parameters are grounded in domain-specific operational standards rather than arbitrary choices, ensuring hydrological relevance. This principle of domain-informed synthetic generation has proven effective in other Earth observation contexts, where geo-typical synthetic data tailored to target region characteristics enables generalization without extensive annotation [69]. Complementarity with operational validation on Canadian data (Section 2.3.2) provides stronger evidence than either approach alone: synthetic evaluation demonstrates learned concepts under controlled conditions, while operational evaluation confirms practical transfer.
Table 4: Comparison of TSAD benchmarks with synthetic or injected anomalies
Benchmark Year Venue Domain # Types Injection Method
Yahoo S5 [40] 2015 Webscope Web KPIs 2 Synthetic + real (outliers, changepoints)
NAB [41] 2015 ICML AI IoT or streaming Real + artificial anomalies
SWaT [46] 2016 CySWater Water treatment 36 Cyberattacks injected into real testbed
WADI [3] 2017 CySWater Water distribution 15 Cyberattacks injected into real testbed
Exathlon [28] 2021 VLDB Distributed systems 6 Disturbances injected into Spark cluster
TODS [38] 2021 NeurIPS D&B General 5 100% synthetic (point, shapelet, seasonal, trend)
TSB-UAD [56] 2022 VLDB General 4 Transformations (noise, smoothing, outliers)
GutenTAG [64] 2022 VLDB General 10 Configurable synthetic generator
TimeEval [85] 2022 VLDB General 10 Toolkit with GutenTAG generator
TSAGen [82] 2022 TNSM AIOps or KPI 3+ Synthetic (season, noise, trend components)
MADE [57] 2024 J Hydrology Hydrology 2 Synthetic pulses and trends
Ours 2025 Hydrology 18 Injection on real USGS data
@@
Mathematical Linear drift, basic offsets Exponential, sigmoid, or polynomial variants (3 to 4 per type) Require functional abstraction
Temporal 8-to-96-hour segments Micro (3 to 58 h), meso (7 to 192 h), macro (72 to 520 h) Probe scale-invariant detection
Parameter Coverage 10-to-32%, nominal severity Coverage trimodal: 3-to-9% or 32-to-44% or 44-to-60%; severity bimodal Test distribution boundaries
Geographic separation: 0 site overlap eliminates memorization of station-specific patterns. Mathematical separation: Each test anomaly type implements 3 to 4 equation variants (eg drift uses linear, exponential, sigmoid, or polynomial forms) that produce similar visual patterns through different mechanisms. Temporal separation: Structured duration regimes (micro, meso, macro) span 3 to 520 hours, with macro durations extended to 520 hours to accommodate documented ice persistence in northern river systems [7]. Parameter separation: Coverage and severity distributions avoid overlap with training ranges, testing generalization at distribution boundaries.
Anomaly taxonomy and literature grounding.
The 18 anomaly types (Table 6) represent failure modes documented in USGS technical guidance and operational practice. Sensor failures (dropout, flatline, spike) reflect data transmission and equipment malfunctions described in USGS electronic processing standards [63]. Hydraulic phenomena (backwater, ice backwater, debris effect, sedimentation) follow stage-discharge relationships documented in WSP 2175 [59] and TWRI 3-A10 [31]. Gradual degradation patterns (drift, rating drift, sensor fouling) align with calibration decay and morphological evolution described by Mansanarez et al [45] and the driftR methodology [67]. Santos-Fernandez et al [61] report empirical prevalence rates from labeled USGS data: drift (4.26%), large spikes (0.13%), small spikes (0.16%), and flatlines or calibration errors (1.09%), confirming these failure modes occur in operational archives.
@@
Category Types Literature Basis Variants
Sensor failures Dropout, flatline, spike USGS electronic standards [63]; Leigh et al [44] 3 to 4 each
Hydraulic phenomena Backwater, ice backwater, debris effect, sedimentation WSP 2175 [59]; TWRI 3-A10 [31] 3 to 4 each
Gradual degradation Drift, rating drift, sensor fouling Mansanarez et al [45]; driftR [67] 3 to 4 each
Processing errors Bias step, desync, quantization USGS time correction standards [63]; Horner et al [26] 3 to 4 each
Complex artifacts Splice, noise burst, gate operation Operational documentation [59] 3 to 4 each
@@
All test anomalies are injected in denormalized physical space to respect discharge-stage coupling. The process involves 3 steps: (1) denormalize to physical units (ft3/s for discharge, ft for stage), (2) apply parametric transformations with hydraulic constraints (eg ice backwater: $H^{\prime}=H(1+\alpha_{\text{ice}})$ and $Q^{\prime}=Q(1-\beta_{\text{ice}})$ with $\alpha\sim U(0.15,0.55)$ and $\beta\sim U(0,0.10)$), and (3) renormalize with clipping. Single-type sequences (30% of the dataset) enable unambiguous per-type evaluation. Compound anomalies (70%, with 40% overlap probability) test discrimination under realistic co-occurrence. Complete injection formulations and reproducibility protocols appear in Appendix F.
Anomaly visualization dashboard.
Figure 2 shows a representative single-segment injected anomaly example with paired clean and corrupted discharge and stage signals, residual diagnostics, rating-curve impact, and site context. For additional qualitative examples spanning all 18 anomaly types and equation form variants, see the anomaly visualization dashboard provided in the HydroGEM Hugging Face repository.
Figure 2: Single-segment synthetic injection example illustrating paired clean versus corrupted discharge and stage time series, the affected interval, residual diagnostics, rating-curve impact, and site context.
@@ 2.3.2 Canadian Zero-Shot Transfer Evaluation
The USGS archive provides quality-controlled products but lacks systematic raw-versus-corrected record pairs that expose operational correction workflows. To test real-world transfer, we evaluate HydroGEM on Environment and Climate Change Canada (ECCC) stations that distribute hourly raw and corrected unit values for both discharge and stage.
@@
Zero-shot protocol: HydroGEM was trained exclusively on USGS sites and never fine-tuned on Canadian data. All experiments are strictly zero-shot, testing whether learned representations generalize across political boundaries, agencies, instrumentation protocols, and rating-curve derivation methods.
Data preprocessing: We obtained ECCC unit value archives, aggregated to hourly resolution, and converted to USGS-aligned imperial units (stage: m to ft via 3.28084; discharge: m3/s to cfs via 35.3147). Temporal alignment ensured all 4 series (stage raw, stage corrected, discharge raw, discharge corrected) were available at each hourly timestep.
@@
Quality filtering: To ensure meaningful evaluation, we applied hydrologically-motivated checks to corrected records: sufficient variability (CV >0.10>0.10), monotonic rating ($\rho>0.5$), reasonable exponent ($b\in[0.5,10]$), moderate fit ($R^{2}\geq 0.3$), valid data fraction ($\geq 70\%$), and limited flatlines ($30\%$ of differences <0.001<0.001 ft). Station-level filtering excluded sites losing $5\%$ of timesteps. Detailed filtering protocols appear in Appendix H.
@@
Weak label construction: We compute relative changes induced by operational correction: $\Delta Q_{\text{rel}}=|Q_{\text{corr}}-Q_{\text{raw}}|/(|Q_{\text{raw}}|+\epsilon)$, and similarly for $H$. Timesteps with $\Delta Q_{\text{rel}}>0.01$ or $\Delta H_{\text{rel}}>0.01$ are marked anomalous ($y_{\text{corr}}=1$), providing weak labels that encode where human experts deemed raw data untrustworthy. The 1% threshold is intentionally conservative, capturing meaningful hydrograph adjustments. These labels indicate where operational corrections occurred rather than providing exhaustive anomaly ground truth. We therefore report tolerant and segment-level metrics in Section 5.
@@
These labels are weak because human editors may miss subtle anomalies or apply corrections for reasons orthogonal to sensor data quality (eg re-rating after channel changes). Nevertheless, they provide externally-validated supervision unavailable in synthetic benchmarks. Pattern-based labels (local correlation deviations) complement correction-based labels by emphasizing multi-point segments where corrected hydrograph shape deviates from raw.
@@
Site sampling: From all stations satisfying quality criteria, we randomly sampled 100 distinct stations for evaluation, ensuring representative out-of-sample behavior without cherry-picking. Window filtering required 10-to-40% correction fraction, identifying substantial but not overwhelming editing effort. For each site, we evaluate HydroGEM in pure zero-shot configuration, comparing predicted anomaly masks against correction-based and pattern-based weak labels.
@@ 3 HydroGEM Model Architecture
Having established the data preparation pipeline in Section 2, we now describe the HydroGEM model that processes the hierarchically normalized 12-dimensional feature vectors to detect anomalies and propose corrections for expert review in streamflow time series (Figure 3). The model input $\mathbf{X}\in\mathbb{R}^{T\times 12}$ consists of the features defined in Section 2.2.1, transformed through the three-tier normalization scheme (Section 2.2.2), and segmented into 576-hour windows (Section 2.1.4). Each sequence represents a continental-scale sample spanning diverse hydrologic regimes, requiring the architecture to learn scale-invariant patterns while respecting physical constraints.
Figure 3: HydroGEM training and inference pipeline. Stage 1 pretrains a hybrid TCN-Transformer backbone (14.2M parameters) using masked reconstruction on clean USGS data. Stage 2 freezes the backbone and trains a lightweight detection head ( 10K parameters) on synthetically corrupted sequences. At inference, the model provides anomaly probabilities with uncertainty estimates via MC Dropout, and suggested corrections for human review.